roulette wheel spins with 38 slots: 18 red,18 black, and 2 green. A ball is spun and lands in a slot. Each slot has an equal chance of capturing the ball. roulette wheel spins 3 consecutive times and a ball lands each time on a red slot. What is the probability that the ball will land on a red slot on the next spin?


1
Expert's answer
2021-09-30T08:06:40-0400

And odds against the ball landing in green slot are 36/38 or 18/19. It can also be found as 1-odds of ball landing in green slot i.e. "1\u20132\/38=>(38\u20132)\/38=>36\/38"

The outcomes of one trial are independent of other trial

"P(red)=\frac{18}{38}\\
P(black)=\frac{18}{38}\\
P(green)=\frac{2}{38}\\
\implies \frac{(\frac{18}{38})^4}{(\frac{18}{38})^3}\\
P= 0.47"